# Inference-time Scaling of Diffusion Models through Classical Search

### v1

@@ 4 Methods
4.1 Local Search via Langevin MCMC with Verifier Gradient
@@
To sample from the composition distribution $\tilde{p}_{0}$, we conduct local-search with hill-climbing methods, aiming to find the local maximum with high $\tilde{p}_{0}$. Specifically, we view the sampling problem as compositional optimization in measure space wibisono2018sampling , and follow the gradient flow of KL-divergence, performing Langevin MCMC steps (details see Appendix. C.1).
@@
Similar to annealed Langevin MCMC in du2023reduce , we could construct a series of annealed functions $\hat{f}_{t}({\bm{x}}_{t})$ with $\hat{f}_{0}({\bm{x}}_{0})=f({\bm{x}}_{0})$. Then we sample from the distributions $\tilde{q}_{t}({\bm{x}}_{t})\propto q_{t}({\bm{x}}_{t})\hat{f}({\bm{x}}_{t})$ through Langevin MCMC in Eq. 2 (details see Appendix. C.2). However, it remains unclear how to design such annealed verifier functions $\hat{f}_{t}({\bm{x}}_{t})$ efficiently, especially without training. Alternatively, training-free guidance in Eq. 3 utilizes the gradient of $f({\bm{x}}_{0|t})$ to optimize ${\bm{x}}_{t}$, which can be computed directly using the diffusion model output. However, naive gradient updates have been observed to produce OOD and adversarial samples shen2024understanding . In ye2024tfg , recurrence (Eq. 4) was found to help avoid adversarial samples in challenging guidance tasks, though its theoretical underpinnings remain poorly understood. We unify these two approaches by demonstrating that training-free guidance with recurrence, in the continuous limit, constitutes an instance of Langevin MCMC. For details see Appendix. C.3.
@@
Thus, the recurrence step (without guidance) can be interpreted as Langevin MCMC applied to the original distribution of the diffusion model $q_{t}({\bm{x}}_{t})$, effectively guiding the samples back toward the data manifold (details see Appendix. C.4). The guidance term then serves as defining a practical annealing path $\hat{f}_{t}({\bm{x}}_{t})$, biasing the sampling path. We are the first to propose this theoretical unification of the two lines of work, providing insights into efficient local search of diffusion models via gradients.
@@
To implement local search, we parameterize the reverse transition kernel $\tilde{p}_{\theta}({\bm{x}}_{t-1}|{\bm{x}}_{t})$ as a sequence of Langevin MCMC steps in Eq. 2, followed by a denoising step with DDIM (Eq. 10) or DDPM (Eq.11). Owing to the equivalence between Langevin MCMC and training-free guidance, we can efficiently tune hyper-parameters using the unified design space of ye2024tfg , with details in Appendix. C.5. We scale up the number of local search steps and observe significant improvements across multiple tasks (Fig. 4(d)), highlighting the effectiveness of local search.
4.2 Global Search for Mode Identification
Figure 2: Illustration of global tree search algorithms. Left: Best-of-N sampling denoises multiple samples in parallel and selects the one with the highest verifier score at the final step. Middle Left: BFS-Resampling resamples particles at intermediate noise levels, allocating more children to nodes with higher scores. Middle Right: BFS-Pruning removes particles with low verifier scores during denoising, sampling at most one child per node. Right: DFS denoises a single particle until its verifier score violates the constraint, then backtracks by adding noise to a previous level and continuing the search.
Although Langevin MCMC is asymptotically exact, it often struggles to transition between distant modes latuszynski2025mcmc due to its inherently local nature. This limitation is especially pronounced in the context of the multimodal distributions induced by diffusion models. To mitigate this, we propose identifying high-score modes via a global search strategy.
We model the Markov chain of the denoising process as a fixed-depth tree, where the transition kernel is given by $\tilde{p}_{\theta}({\bm{x}}_{t-1}|{\bm{x}}_{t})$. As discussed earlier, when applying local search, we parameterize the transition kernel with multiple Langevin MCMC steps, and otherwise we directly adopt conventional diffusion samplers song2020denoising ; ho2020denoising ; lu2022dpm . Hence, our perspective unifies both stochastic and deterministic sampling with guided sampling. This abstraction decouples the transition kernel from the sampling process, enabling the application of classic tree search heuristics to develop compute-efficient exploration methods.
BFS-style linear search
Inspired by breadth-first search (BFS), which expands nodes level-by-level, we denoise a set of particles in parallel at each noise level. The simplest approach is best-of-$N$ sampling: generate $N$ candidate trajectories and select the one with the highest verifier score at the final step. While simple, this strategy ignores information from intermediate stages.
To improve efficiency, similar to best-first search pearl1984heuristics , we evaluate each intermediate particle $\left\{{\bm{x}}_{t}^{k}\right\}_{k=1}^{K_{t}}$ using its verifier score $f({\bm{x}}_{0|t}^{k})$, where $K_{t}$ is the number of particles at timestep $t$. Based on these scores, we dynamically reallocate computation by sampling more children for high-scoring nodes via the reverse transition kernel $\tilde{p}_{\theta}({\bm{x}}_{t-1}|{\bm{x}}_{t})$. This forms the basis of BFS-Resampling, where each node samples children in proportion to its normalized verifier score
$n_{t}^{k}=\texttt{Round}\left(K_{t}\frac{f({\bm{x}}_{0|t}^{k})^{\tau_{t}}}{\sum_{i=1}^{K_{t}}f({\bm{x}}_{0|t}^{i})^{\tau_{t}}}\right)\,.$
We then sample $n_{t}^{k}$ children from ${\bm{x}}_{t}^{k}$ using $\tilde{p}_{\theta}({\bm{x}}_{t-1}|{\bm{x}}_{t})$ and proceed to the next denoising step.
When using deterministic ODE solvers lu2022dpm as transition kernels, it is ineffective to sample multiple children from the same parent. To address this, we propose BFS-Pruning, a variant that retains only top-scoring nodes and allows at most one child per parent $n_{t}^{k}=\min\left(1,n_{t}^{k}\right)$. This leads to a gradual narrowing of the search tree as $K_{t}$ decreases over time, enabling us to explore more diverse modes with more initial particles under the same compute.
As shown in Fig. 6, both resampling and pruning significantly improve efficiency over standard best-of-$N$ sampling. The detailed pseudo code is in Alg. 3.
DFS-style non-linear search
Depth-first search (DFS) explores one branch of the search tree as deeply as possible before backtracking. In our setting, this corresponds to iteratively denoising a single particle until its verifier score drops below a predefined threshold: $f({\bm{x}}_{0|t})\leq\delta_{t}$, where $\delta_{t}$ is a scheduled threshold for timestep $t$.
Once the constraint is violated, the algorithm backtracks by reintroducing noise, moving to a higher noise level $t_{\text{next}}=t+\Delta_{T}$ using the forward diffusion process $q({\bm{x}}_{t_{\text{next}}}|{\bm{x}}_{t})$ in Eq. 1. This allows the model to restart the denoising process from a different region of the manifold, encouraging exploration of diverse modes. By placing a cap on the number of backtracking steps, we can explore the space globally while keeping computation bounded.
A key strength of DFS is its ability to allocate compute adaptively: difficult prompts and low-quality trajectories naturally trigger more backtracking and exploration, while easier instances are solved more directly. This dynamic behavior is driven purely by the verifier signal, without needing to know the difficulty in advance. As shown in Fig. 5, this strategy leads to substantial gains in efficiency and performance. The detailed pseudo code is in Alg. 4.
@@ 5 Experiments
In this section, we apply inference-time scaling with our search strategy across a range of domains. In Sec.5.1, we show that diffusion models can generate long-horizon plans in complex maze environments with inference-time scaling, given a world model as verifier. In Sec.5.2, we leverage a pretrained Q-function as a verifier to guide diffusion policies, achieving strong performance on D4RL fu2020d4rl locomotion benchmarks. Finally, in Sec. 5.3, we approach the image generation task, and demonstrate that our global search algorithms consistently outperform best-of-N sampling in terms of efficiency, and we reveal interesting properties of BFS and DFS.
5.1 Long Horizon Planning
Figure 4: (Illustration and results for maze planning) Fig. 4(a) shows the maze layout of Giant (left) and Ultra (right), with start: $\bullet$ and goal: $\star$. Fig. 4(b) illustrates the effect of local search, where left shows the directly sampled trajectory that violates the maze block on the bottom left, and right shows the successful sample after local search. Fig. 4(c) shows that inference time search generates diverse solutions. Fig. 4(d) shows the performance with scaling number of local search steps with Langevin MCMC. Fig. 4(e) and Fig. 4(f) shows the scaling behaviors of the global search methods for PointMaze Giant and PointMaze Ultra respectively. Details see Appendix. E.1.
Diffusion models have been widely adopted in planning for trajectory synthesis ubukata2024diffusion . We evaluate long-horizon planning in the PointMaze environment from OGBench park2024ogbench , training the diffusion model following the approach in Diffuser janner2022planning , where the full trajectory is denoised:
To highlight the effectiveness of inference-time scaling, we evaluate on the largest PointMaze Giant environment and introduce an even more challenging variant, PointMaze Ultra, following the protocol from marcucci2023motion . The maze layouts are shown in Fig. 4(a). We define the verifier loss for each colliding point of the trajectory as the squared distance to its nearest wall, thus the overall verifier score measures the degree the trajectory violates the environment (details see Appendix. E.1). This verifier functions as a physics-grounded instance of a world model. Importantly, naively maximizing the verifier score does not guarantee a successful plan, and planning remains difficult even with full access to the maze layout luo2024potential ; marcucci2023motion .
In the most challenging scenario—navigating from the bottom-left to the top-right corner—naive sampling strategies rarely succeed, necessitating multiple local search steps (Fig. 4(d)). As illustrated in Fig. 4(b), local search guided by the verifier enables the model to find successful plans that navigate around obstacles. As shown in Fig. 4(e) and Fig. 4(f), incorporating global search alongside local refinement leads to over 90% success rates in both maze environments, with both BFS and DFS significantly improving search efficiency compared to naive best-of-$N$ sampling. Moreover, as shown in Fig. 4(c), global search enables the generation of diverse, high-quality solutions.
5.2 Offline Reinforcement Learning
Table 1: Performance on D4RL locomotion tasks, score normalized as suggested in fu2020d4rl . Following lu2023contrastive we report the mean and standard deviation over 5 independent trials, highlighting the numbers within 5 percent of the maximum in every individual task. To faithfully report our performance, we conduct the hyper-parameter search and evaluation with different seeds. For more details see Appendix. E.2.
Dataset Environment CQL BCQ IQL SfBC DD Diffuser D-QL QGPO TTS(ours)
Medium-Expert HalfCheetah 62.4 64.7 86.7 92.6 90.6 79.8 96.1 93.5 $\mathbf{93.9\pm 0.3}$
Medium-Expert Hopper 98.7 100.9 91.5 108.6 111.8 107.2 110.7 108.0 $104.4\pm 3.1$
Medium-Expert Walker2d 111.0 57.5 109.6 109.8 108.8 108.4 109.7 110.7 $\mathbf{111.4\pm 0.1}$
Medium HalfCheetah 44.4 40.7 47.4 45.9 49.1 44.2 50.6 54.1 $\mathbf{54.8\pm 0.1}$
Medium Hopper 58.0 54.5 66.3 57.1 79.3 58.5 82.4 98.0 $\mathbf{99.5\pm 1.7}$
Medium Walker2d 79.2 53.1 78.3 77.9 82.5 79.7 85.1 86.0 $\mathbf{86.5\pm 0.2}$
Medium-Replay HalfCheetah 46.2 38.2 44.2 37.1 39.3 42.2 47.5 47.6 $\mathbf{47.8\pm 0.4}$
Medium-Replay Hopper 48.6 33.1 94.7 86.2 100.0 96.8 100.7 96.9 $\mathbf{97.4\pm 4.0}$
Medium-Replay Walker2d 26.7 15.0 73.9 65.1 75.0 61.2 94.3 84.4 $79.3\pm 9.7$
Average (Locomotion) 63.9 51.9 76.9 75.6 81.8 75.3 86.3 86.6 86.1
Recently, diffusion models emerged as a powerful action prior in robotics due to its ability to model complex and multimodal distributions chi2023diffusion ; liu2024rdt . However, these diffusion policies are typically trained on offline datasets and struggle to adapt to reinforcement learning or test-time requirements. Following prior work peters2010relative , we formulate the offline RL problem as sampling from a Q-regularized distribution:
$\pi^{*}(\bm{a}|\bm{s})\propto\mu(\bm{a}|\bm{s})e^{\beta Q_{\psi}(\bm{s},\bm{a})}\,,$
where $Q_{\psi}$ is a learned Q-function representing preferences over actions, and $\mu$ is the behavior policy, which we model using a diffusion prior.
In principle, exact sampling from $\pi^{*}$ using diffusion models requires training a noise-dependent energy function on noisy actions $\bm{a}_{t}$ janner2022planning ; lu2023contrastive . However, this approach is often impractical in real-world settings. First, it requires evaluating the pretrained Q-function on samples from the diffusion model, which is costly due to the difficulty of data collection in robotics. Second, many Q-functions rely on foundation models as backbones brohan2023can ; nakamotosteering , and training them on noisy samples risks degrading their pretrained knowledge. In contrast, test-time search (TTS) offers a promising alternative. By decoupling policy and value, TTS enables flexible composition with foundation models without retraining.
We evaluate TTS using both local and global search strategies on the D4RL locomotion benchmarks fu2020d4rl . Following the setup in lu2023contrastive , we adopt their pretrained diffusion policy and ground-truth Q-function. Results are presented in Table 1, where our approach achieves performance competitive with the best training-based methods, including QGPO lu2023contrastive and Diffusion-QL wang2022diffusion . In comparison, our approach is training-free, requiring only an off the shelf Q function and diffusion model.
5.3 Image Generation
In this section we evaluate inference-time scaling in image generation tasks, especially the efficiency of global search.
5.3.1 Compositional Text-to-Image Generation
Figure 5: CompBench huang2023t2icompbench text-to-image results. Fig. 5(a) and Fig. 5(b) show DFS performance on attribute binding and object relationship tasks, achieving up to 2× lower computational cost than Best-of-N. Fig. 5(c) presents the average compute allocated by DFS for prompts of increasing difficulty, evaluated on the color dataset.
We evaluate inference-time scaling for text-to-image generation on the CompBench dataset huang2023t2icompbench using the SSD-1B model gupta2024progressive . For the search verifiers, we use task-specific oracle verifiers: a BLIP-VQA model li2022blip for the attribute binding task and a UniDet model zhou2022simple for the object relationship task. Since these verifiers are non-differentiable, we apply global search exclusively. Detailed setup is in Appendix. E.3.1.
Improved efficiency and adaptivity with DFS
We evaluate Best-of-N and DFS on the five datasets in attribute binding and object relationship tasks, with each dataset consists of 300 prompts. The results are shown in Fig. 5(a) and Fig. 5(b), where DFS consistently outperforms Best-of-N while using up to 2× less compute. To evaluate DFS’s ability to adaptively allocate computation, we divide the 300 prompts in the color dataset into four difficulty bins based on their average score over 8 samples. Using the same set of hyperparameters across all prompts, Fig. 5(c) shows that DFS allocates less compute to easier prompts (bin 1) and progressively more to harder ones with lower scores—demonstrating automatic, difficulty-aware compute allocation. Unlike snell2024scaling , which manually selects different search strategies based on predefined difficulty labels, our method requires no prior knowledge of difficulty. DFS dynamically adjusts compute allocation during inference using a single, unified algorithm and fixed parameters.
5.3.2 Conditional ImageNet Generation
Conditional generation of diffusion models using a unconditional foundation model has important applications ye2024tfg as foundation models obtain increasing expressivity. However, classifier guidance dhariwal2021diffusion requires training a noise-dependent classifier on samples generated by the diffusion model, which is difficult since it may require a lot of attempts for an unconditional model to generate the correct class. In this section, we treat the pretrained classifier as the verifier function and conduct inference-time scaling with both local and global search. The experimental details are in Appendix. E.3.2.
Table 2: Best-of-N results for ImageNet conditional generation, with FID and Accuracy averaged across the two labels.
#Particles Single verifier Double-verifier
@@ 5.3.2 Conditional ImageNet Generation
2 188.5 25.6% 178.4 29.7%
4 171.5 31.8% 151.2 37.5%
8 155.7 35.8% 127.8 49.2%
Mitigating verifier hacking with double-verifier
One of the major challenges of guiding the samples using the verifier gradient is that image classifiers are prone to adversarial examples szegedy2013intriguing . Inspired by the double-Q learning in RL van2016deep , we use two verifiers for local search gradient computation and global search respectively. Specifically, we use two different vision transformers xu2023imagereward with input resolution 224 and 384. For accuracy evaluation we use another classifier from touvron2021training . As shown in Table. 2, using double-verifier consistently improves performance with Best-of-N sampling, with up to 2x less compute.
Figure 6: Results for ImageNet conditional generation with BFS Fig. 6(a) and Fig. 6(b) show the FID and class accuracy of conditional samples. BFS consistently outperform Best-of-N sampling in efficiency.
Improved Efficiency with BFS
We investigate breadth-first search (BFS) as a means to improve the computational efficiency of Best-of-N sampling. All experiments use the double-verifier setup, where one verifier guides gradient-based local search and the other handles resampling or pruning. As shown in Fig. 6, both BFS-Resampling and BFS-Pruning outperform Best-of-N in terms of compute efficiency. BFS-Pruning is more effective under limited compute budgets, while BFS-Resampling achieves higher performance with increased compute.

### v2

@@ 4.1.2 DFS-style non-linear search
4.2 Scaling Local Search via Langevin MCMC with Verifier Gradient
@@
Global search can efficiently discover the high score modes from the base diffusion model, but can not generate higher quality samples that exceed the pretrained model. Thus, we aim to sample from the compositional distribution $\tilde{p}_{0}$ in Eq. 5 for higher quality samples. To optimize the compositional objective, we conduct local-search with hill-climbing methods, aiming to find the local maximum with high $\tilde{p}_{0}$. Specifically, we view the sampling problem as compositional optimization in measure space (Wibisono, 2018), and follow the gradient flow of KL-divergence, performing Langevin MCMC steps (details see Appendix. C.1).
@@
Similar to annealed Langevin MCMC in Du et al. (2023), we could construct a series of annealed functions $\hat{f}_{t}({\bm{x}}_{t})$ with $\hat{f}_{0}({\bm{x}}_{0})=f({\bm{x}}_{0})$. Then we sample from the distributions $\tilde{q}_{t}({\bm{x}}_{t})\propto q_{t}({\bm{x}}_{t})\hat{f}({\bm{x}}_{t})$ through Langevin MCMC in Eq. 2 (details see Appendix. C.2). Alternatively, training-free guidance in Eq. 3 utilizes the gradient of $f({\bm{x}}_{0|t})$ to optimize ${\bm{x}}_{t}$, which can be computed directly using the diffusion model output. However, naive gradient updates have been observed to produce OOD and adversarial samples (Shen et al., 2024). In Ye et al. (2024), recurrence (Eq. 4) was found to help avoid adversarial samples in challenging guidance tasks, though its theoretical underpinnings remain poorly understood. We unify these two approaches by demonstrating that training-free guidance with recurrence, in the continuous limit, constitutes an instance of Langevin MCMC. For details see Appendix. C.3, and a rigorous convergence bound is in Theorem. 1.
@@
Thus, the recurrence step (without guidance) can be interpreted as Langevin MCMC applied to the original distribution of the diffusion model $q_{t}({\bm{x}}_{t})$, and the guidance term $\bm{\Delta}_{t}$ in Eq. 3 then serves as defining a practical annealing path $\hat{f}_{t}({\bm{x}}_{t})$ that bias the sampling path towards high reward areas beyond the modes of the base model. We are the first to propose this theoretical unification of the two lines of work, providing insights into efficient local search of diffusion models via gradients.
@@
We implement local search by parameterizing the reverse transition kernel $\tilde{p}_{\theta}({\bm{x}}_{t-1}|{\bm{x}}_{t})$ as a sequence of Langevin MCMC steps (Eq. 2), followed by a denoising step using DDIM (Eq. 11) or DDPM (Eq. 12); see Appendix C.5 for details. Unlike classifier-guidance or naive training-free guidance, which apply only gradient guidance in the denoising step, our approach incorporates explicit Langevin MCMC steps. In Sec. 5.3, we scale the number of local search steps for the first time and observe substantial improvements over pretrained models across multiple tasks.
@@ 5 Experiments
In this section, we apply inference-time scaling with our search strategy across a range of domains. In Sec. 5.1, we present a strengthened BFS baseline that outperforms previous particle-based methods. In Sec. 5.2, we demonstrate the adaptivity and efficiency of our DFS method. In Sec. 5.3, we scale up local search in challenging decision-making domains, highlighting the importance of jointly scaling local and global search. In Sec. 5.4, we propose a double-verifier strategy to mitigate the reward hacking problem in inference-scaling.
5.1 Elucidating the Design Space of BFS
In this section, we explore the design choices of BFS, and present an improved BFS (Increase, Max, SSP) over prior particle-based or tree-search-based methods (Singhal et al., 2025b; Kim et al., 2025; Guo et al., 2025; Li et al., 2024). To ensure a fair comparison, we directly use the official implementation of FK-steering (Singhal et al., 2025a) with the ImageReward (Xu et al., 2023a) verifier and the SD v1.5 model. For details, see Appendix E.1.
Table 1: Ablation of BFS design choices. For each section, we vary one design choice while keeping the others fixed. Left: Resampling methods (with Constant tempering, Max scoring). Center: Scoring functions (with SSP resampling, Constant tempering). Right: Tempering schedules (with SSP resampling, Max scoring).
Resampling Scoring Tempering
N BoN Multinomial SSP Current Difference Max Increase Inf Constant
4 $0.702\pm 0.057$ $0.743\pm 0.037$ $\mathbf{0.834\pm 0.041}$ $0.812\pm 0.037$ $0.823\pm 0.036$ $\mathbf{0.834\pm 0.041}$ $\mathbf{0.882\pm 0.029}$ $0.667\pm 0.076$ $0.834\pm 0.041$
8 $0.896\pm 0.031$ $0.926\pm 0.042$ $\mathbf{1.032\pm 0.035}$ $0.996\pm 0.029$ $1.013\pm 0.032$ $\mathbf{1.032\pm 0.035}$ $\mathbf{1.087\pm 0.031}$ $0.775\pm 0.087$ $1.032\pm 0.035$
As mentioned in Sec.4.1.1, we experiment with different resampling, scoring, and tempering design choices. We begin with BFS (Constant, Max, Multinomial) and evaluate various resampling strategies. As shown in Table 1 (left), SSP significantly outperforms naive multinomial resampling, and we adopt it in our design. We then assess scoring functions and tempering schedules under SSP resampling, as reported in Table 1 (center) and 1 (right). Here, we observe modest improvements with Max scoring and Increase tempering, leading to our final design: BFS (Increase, Max, SSP).
Model N BoN FKSinghal et al. (2025a) DASKim et al. (2025) TreeG Guo et al. (2025) SVDDLi et al. (2024) BFS (ours)
SD v1.5 4 $0.702\pm 0.057$ $0.743\pm 0.037$ $0.878\pm 0.028$ $0.860\pm 0.033$ $0.667\pm 0.076$ $\mathbf{0.882\pm 0.029}$
SD v1.5 8 $0.896\pm 0.031$ $0.926\pm 0.042$ $1.052\pm 0.033$ $1.023\pm 0.018$ $0.775\pm 0.087$ $\mathbf{1.087\pm 0.031}$
@@ 5.1 Elucidating the Design Space of BFS
SDXL 8 $1.198\pm 0.021$ $1.251\pm 0.011$ $1.265\pm 0.019$ $1.261\pm 0.021$ $1.225\pm 0.027$ $\mathbf{1.291\pm 0.018}$
Table 2: Comparison of our BFS with prior methods
To compare our improved BFS (Increase, Max, SSP) with prior baselines, we additionally experiment with the SDXL model (Podell et al., 2023), which differs from the model used in our ablations. As shown in Table 2, our improved BFS consistently outperforms previous methods across compute budgets and models. In the following experiments, we use the improved BFS as our baseline.
5.2 Adaptive and Efficient Inference-Scaling with DFS
Figure 3: CompBench text-to-image results with DFS. Left: the Pareto curve of DFS, with DFS-$\delta$ denotes DFS with threshold $\delta_{t}=\delta$. Right: average compute allocation by DFS for prompts of increasing difficulty.
In this section, we evaluate the adaptivity and efficiency of DFS on the CompBench dataset (Huang et al., 2023), using the SSD-1B model (Gupta et al., 2024) and a VLM (Li et al., 2022) as our verifier. The detailed setup is provided in Appendix E.2. Through these experiments, we address the following questions:
  • Can DFS outperform Best-of-N and prior particle-based methods? As shown in Fig. E.2 (left), DFS consistently outperforms Best-of-N and also our improved BFS design across different compute budgets, with up to $2\times$ less computational cost.
  • Can DFS adjust compute allocation with different thresholds? We evaluate DFS across a wide range of practical threshold values ($0.5$, $0.7$, and $0.9$) and find that lower thresholds automatically allocate less compute, while higher thresholds scale up compute for better quality. DFS consistently outperforms baseline methods across all threshold choices, demonstrating the robustness of our method for different hyperparameters.
  • Can DFS dynamically adjust compute allocation for different instances? We measure the computational cost of DFS on prompts of varying difficulty under fixed thresholds, where the difficulty of a prompt is defined as the average score over four random samples. As shown in Fig. 3 (right), harder prompts with lower scores naturally consume more compute, without requiring prior knowledge of difficulty as in Snell et al. (2024).
Unlike linear-search methods that use a fixed exploration schedule, DFS offers higher efficiency and adaptivity, which may be of independent interest to the broader community.
5.3 Joint Scaling Local and Global Search
Although global search methods such as BFS and DFS can efficiently explore the generative space of the diffusion model, they are restricted to the modes of the base distribution and therefore cannot exceed the capabilities of the base model. To optimize the compositional objective in Eq. 5 and sample from high-reward regions beyond the base model, we propose scaling up local search steps via annealed Langevin MCMC, introducing a new scaling dimension for diffusion models. We validate the effectiveness of scaling local search in challenging decision-making domains, such as long-horizon planning and offline RL.
Baselines
To evaluate the effectiveness of scaling local search steps, we compare our method with particle-based DAS (Kim et al., 2025), which also leverages verifier gradients but only incorporates gradient guidance without recurrent local search. We also compare against the state-of-the-art training-free guidance method TFG (Ye et al., 2024), which performs multiple recurrence steps but lacks any global search. For fairness, compute is measured as the total NFEs of both local and global search. As shown in the following experiments, scaling local and global search separately yields suboptimal performance, whereas our joint scaling strategy establishes a new Pareto frontier.
5.3.1 Long Horizon Planning
Figure 4: Pareto curves of local search. Left: Pareto curves of best-of-N with different local search steps, where BoN-$i$ denotes $i$ local search steps. Right: Pareto curves of BFS and DFS with 6 local search steps.
Figure 5: Illustration of Maze layout and task, with failed trajectory (left) and successfual sample (right).
Diffusion models have been widely adopted in planning for trajectory synthesis (Ubukata et al., 2024). We evaluate long-horizon planning in a challenging PointMaze environment, using the base model trained following Diffuser (Janner et al., 2022), with the verifier defined as the total number of collisions between the trajectory and maze walls (see Appendix E.3 for details). (Luo et al., 2024; Marcucci et al., 2023). As shown in Fig. 5 (left), naive sampling without local search often resulted in failed trajectories that violates the maze layout in some of the segments. When scaling up local search steps as in Fig. 5 (right), we are able to generate successful plans free of violation.
To evaluate the compute efficiency of recurrent steps, we observe in Fig. 4 (left) that scaling up local search with BoN improves the overall Pareto frontier and significantly outperforms baseline methods. Scaling local search only in TFG (Ye et al., 2024) is efficient with a low compute budget but fails to scale with increased compute, as local search alone can become trapped in local optima. DAS (Kim et al., 2025) is more efficient than the corresponding BoN-0 baseline without local search, but underperforms best-of-N when more local search steps are used. In Fig. 4 (right), we show that local search can be combined with global search techniques such as BFS and DFS to further improve scaling efficiency, demonstrating the flexibility of our framework.
5.3.2 Offline Reinforcement Learning
Table 3: Performance on D4RL locomotion tasks, highlighting results within 5% of the maximum.
Dataset Environment IQL SfBC $D$ Diffuser D-QL QGPO TFG DAS TTS(ours)
Medium-Expert HalfCheetah 86.7 92.6 90.6 79.8 96.1 93.5 $90.2\pm 0.2$ $\mathbf{93.3\pm 0.3}$ $\mathbf{93.9\pm 0.3}$
Medium-Expert Hopper 91.5 108.6 111.8 107.2 110.7 108.0 $100.2\pm 3.5$ $105.4\pm 5.1$ $104.4\pm 3.1$
Medium-Expert Walker2d 109.6 109.8 108.8 108.4 109.7 110.7 $\mathbf{108.1\pm 0.1}$ $\mathbf{111.4\pm 0.1}$ $\mathbf{111.4\pm 0.1}$
Medium HalfCheetah 47.4 45.9 49.1 44.2 50.6 54.1 $\mathbf{53.1\pm 0.1}$ $\mathbf{53.4\pm 0.1}$ $\mathbf{54.8\pm 0.1}$
Medium Hopper 66.3 57.1 79.3 58.5 82.4 98.0 $\mathbf{96.2\pm 0.5}$ $71.3\pm 2.7$ $\mathbf{99.5\pm 1.7}$
Medium Walker2d 78.3 77.9 82.5 79.7 85.1 86.0 $\mathbf{83.2\pm 1.4}$ $\mathbf{83.9\pm 0.9}$ $\mathbf{86.5\pm 0.2}$
Medium-Replay HalfCheetah 44.2 37.1 39.3 42.2 47.5 47.6 $\mathbf{45.0\pm 0.3}$ $42.2\pm 0.1$ $\mathbf{47.8\pm 0.4}$
Medium-Replay Hopper 94.7 86.2 100.0 96.8 100.7 96.9 $93.1\pm 0.1$ $\mathbf{96.7\pm 3.0}$ $\mathbf{97.4\pm 4.0}$
Medium-Replay Walker2d 73.9 65.1 75.0 61.2 94.3 84.4 $69.8\pm 4.0$ $63.8\pm 2.0$ $79.3\pm 9.7$
Average (Locomotion) 76.9 75.6 81.8 75.3 86.3 86.6 82.1 80.2 86.1
Recently, diffusion models have emerged as a powerful action prior in robotics due to their ability to model complex and multimodal distributions (Chi et al., 2023). However, these diffusion policies are typically trained on offline datasets and struggle to adapt to reinforcement learning or test-time requirements. Following Peters et al. (2010), we formulate the offline RL problem as sampling from a Q-regularized distribution: $\pi^{*}(\bm{a}|\bm{s})\propto\mu(\bm{a}|\bm{s})e^{\beta Q_{\psi}(\bm{s},\bm{a})}$, where $Q_{\psi}$ is a learned Q-function representing preferences over actions, and $\mu$ is the behavior policy, which we model using a diffusion prior. We approach this problem from the inference-scaling perspective, composing off-the-shelf pretrained diffusion policy with ground-truth Q-functions without training.
Among the baselines, Diffuser (Janner et al., 2022), QGPO (Lu et al., 2023), and D-QL (Wang et al., 2022) require additional joint training of the diffusion model and Q-function. To demonstrate the effectiveness of our method test-time search (TTS), we allow TFG (Ye et al., 2024) and DAS (Kim et al., 2025) to use up to twice the compute of TTS. As shown in Table 3, TTS achieves performance comparable to training-based baselines, while DAS and TFG struggles on the Medium and Medium-Replay datasets where the model’s capabilities are limited. Details see Sec. E.4.
Environment TTS-distill DPPO
Halfcheetah $\mathbf{51.6\pm 0.7}$ 47.8
Hopper $\mathbf{98.8\pm 0.8}$ 92.8
Walker2d $\mathbf{86.3\pm 0.5}$ 82.7
Table 4: Results for policy distillation with TTS
Advancing Online Learning with Policy Distillation
To enhance the capabilities of the base model, we fine-tune it on actions generated by TTS and subsequently evaluate its one-shot sampling performance. We adapt the Medium datasets by replacing the original actions with those produced by TTS, and fine-tune the corresponding models from Lu et al. (2023) (see Sec. E.4.1 for details). Perhaps surprisingly, this simple offline distillation outperforms SOTA diffusion-based online RL methods such as DPPO (Ren et al., 2024), as shown in Table 4. To the best of our knowledge, this is the first work to demonstrate expert iteration with inference scaling for diffusion models, a direction that may be of independent interest.
5.4 Mitigating reward hacking with double verifier
In this section, we show that reward hacking and over-optimization can be mitigated by a double-verifier strategy: employing separate verifiers for local and global search. As observed in Shen et al. (2024), training-free guidance with verifier gradients is vulnerable to adversarial reward hacking: gradient guidance often over-optimize the verifier, causing it to classify them as belonging to the target class despite being out-of-distribution (OOD). Inspired by double-Q learning in reinforcement learning (Van Hasselt et al., 2016), we propose a double-verifier approach, assigning distinct verifiers to local and global search to efficiently reduce overestimation.
Table 5: Results for ImageNet conditional generation with double-verifier using BoN and BFS.
BoN-single BoN-double BFS-single BFS-double
[…]
4 171.5 31.8% 0.161 151.2 37.5% 0.164 156.2 36.1% 0.169 145.5 44.3% 0.181
8 155.7 35.8% 0.165 127.8 49.2% 0.184 133.3 46.5% 0.183 118.2 55.9% 0.214
We evaluate the proposed double-verifier on the challenging conditional ImageNet generation task, generating target-class samples from an unconditional model guided by a pretrained classifier defined only on clean samples. We report the FID and class accuracy of 256 samples. We also report the MSP score (Hendrycks & Gimpel, 2016), with higher MSP score indicating less OOD samples. As shown in Table  5, using double-verifier significantly improves performance over single verifier with 2x less compute. Also, double-verifier significantly reduces OOD samples indicated by the higher MSP score, improving robustness alongside efficiency. Visualizations and details see Appendix. E.5.